OpenStudy (alexandervonhumboldt2):

Prove that if a/b=c/d a, b, c, d are positive numbers and a is the biggest prove that a+d>b+c.

OpenStudy (alexandervonhumboldt2):

@TheSmartOne @Loser66 @ash2326 @bibby @Ashleyisakitty

OpenStudy (alexandervonhumboldt2):

@TuringTest @Directrix

OpenStudy (alexandervonhumboldt2):

@freckles

OpenStudy (freckles):

so we have the following given: \[\frac{a}{b}=\frac{c}{d}\] \[a>b,c,d\] \[a,b,c,d>0\]

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (freckles):

well from the first equation we have \[ad=cb\] ... \[ad+dd=cb+dd\] \[d(a+d)=cb+dd\] \[d(a+d)+bb=cb+dd+bb\] \[d(a+b)+bb=b(c+b)+dd\] hmm... maybe we can think something along this idea Like maybe we can alter this idea or something... what do you think?

OpenStudy (freckles):

oops I changed d to b in the last line there*

OpenStudy (alexandervonhumboldt2):

wel let me think

OpenStudy (alexandervonhumboldt2):

and what then? seems a start is correct

OpenStudy (freckles):

you know ad=bc right? and you know a>b so aa>bb

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (freckles):

so what sign goes in between ad+aa ___ bc+bb

OpenStudy (alexandervonhumboldt2):

>

OpenStudy (freckles):

so we have ad+aa>bc+bb or a(d+a)>b(c+b)

OpenStudy (freckles):

we are given a>b do you think you can say anything else about the inequality a(d+a)>b(c+d) given that a>b

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

need to remove a and b from sides dont know how because they are firrent

OpenStudy (mathmath333):

by contradiction if we assume that this case \(\large \begin{align} \color{black}{a<(b=c=d)\hspace{.33em}\\~\\ }\end{align}\) then \(\large \begin{align} \color{black}{a+d>b+c\hspace{.33em}\\~\\ a+d>d+c\hspace{.33em}\\~\\ a>c\hspace{.33em}\\~\\ }\end{align}\) which contradicts our idea

OpenStudy (alexandervonhumboldt2):

but b c d are not equal

OpenStudy (mathmath333):

yes its only for that case

OpenStudy (alexandervonhumboldt2):

but for all cases?

OpenStudy (alexandervonhumboldt2):

@freckles we are nearly done but dont know last step

OpenStudy (alexandervonhumboldt2):

a*(d+a)=b*(c+b) if we remove a and b we get what we need

OpenStudy (alexandervonhumboldt2):

woops >

OpenStudy (mathmath333):

lol my soln seems wrong since a/b=c/d given

OpenStudy (freckles):

right i was thinking well if we have a/2=2/2 then a is also 2

OpenStudy (freckles):

so they are all equal is b=c=d

OpenStudy (freckles):

if b=c=d*

OpenStudy (alexandervonhumboldt2):

well 6/3=2/1

OpenStudy (alexandervonhumboldt2):

gono offline of a second

OpenStudy (alexandervonhumboldt2):

yeah back

OpenStudy (anonymous):